OpenStudy (anonymous):

Help.. Can some one teach me how to do step by step? Find the equation of the line with Slope = 310 passing through the point (-10,-11). Write your answer in the form y=mx+b.

OpenStudy (anonymous):

m = slope, then you have y = 310x + b. Now you have to plug in x and y to solve for b. Know what I mean?

OpenStudy (anonymous):

no idea what you mean :( sorry

OpenStudy (anonymous):

how do i find x & y?

OpenStudy (anonymous):

That's alright. You already have them thats the coordinates (-10, -11).

OpenStudy (anonymous):

Ok take a random point (x,y) and use your given points. We know that gradient = \[(y_2-y_2)/(x_2-x_1)\] So do (y-11)/(x-(-10))=310 and then cross multiply

OpenStudy (anonymous):

sorry meant \[y_2 - y_1\]

OpenStudy (anonymous):

Ignore beeqay what I said was correct.

OpenStudy (anonymous):

its 3/10 not 310.. sorry

OpenStudy (anonymous):

I will list it in steps. 1. m=(3/10) so rewrite the equation y=mx+b as y=(3/10)x+b 2. Plug in the set of coordinates they give you in this case it is (-10,-11) so the equation turns into -11=(3/10)(-10)+b 3. Solve for b 4. rewrite the equation as y = 3/10x + (whatever number you find for b in step 3) Understand? That's is alright the steps are the same.

OpenStudy (anonymous):

That's alright**

OpenStudy (anonymous):

lol, you don't do what you have been told above if you are given points.

OpenStudy (anonymous):

Beeqay what are you talking about? Read the question again. They need the equation in y=mx+b form. You have the slope and coordinates, but you don't have "b". So obviously you need to solve for "b". Ignore beeqay what I said is correct.

OpenStudy (anonymous):

i got y=.3+-8

OpenStudy (anonymous):

Okay you are close but you don't write your final answer like that. You write y=-3x-8 know what I mean?

OpenStudy (anonymous):

oh , okay :) Thank you

OpenStudy (anonymous):

No problem =)

OpenStudy (anonymous):

haha b is a constant lad, in this case b would be 14 and your equation would be y= (3/10)x+14 m is 3/10 and b=14 But if you do not wish to understand this then I will not force you to.

OpenStudy (anonymous):

Haha man you really don't know when to quit beeqay. I am aware b is a constant and that is what we were solving for because we didn't have that information. If you plug -10 into the equation you wrote you get y=11 and the coordinates are (-10,-11) so clearly you are wrong lad. You are the one who does not understand this obviously.

OpenStudy (anonymous):

Let me know if you get y=-3x-8 correct.

OpenStudy (anonymous):

oh shoot I meant to write y = (3/10)x-8

OpenStudy (anonymous):

Oh I hadn't seen the -10...but y = (3/10)x-8 is correct and you can use whatever method. Because I get the same value if I use -10.

OpenStudy (anonymous):

i put in y=.3x-8 and it was correct /: so im not sure lol im sure both of you are right :)

OpenStudy (anonymous):

Yea that's correct.

OpenStudy (anonymous):

No just me haha, but oh I see you have a decimal next to the 3 yeah thats why because .3=3/10 I thought you just wrote -3 and yeah I just forgot what the slope was, but what beeqay was doing had nothing to do with the question being asked.

OpenStudy (anonymous):

Haha you have a very limited way of thinking lad. So how then did I find the right answer using my method?

OpenStudy (anonymous):

Okay I see how you could solve for b using your method but it is clearly more difficult and if I have a limited way of thinking then so do you because you also stated that my method was incorrect . Even though it is clearly correct and faster.

OpenStudy (anonymous):

Because your answer was incorrect initially. But anyway, the reason why I suggested my method was because at times you are not given the gradient and you have to use the change in y/ change in x so it's good to know the method. But either way, neither of the methods used is hard. I call a truce, I'm tired of arguing.

OpenStudy (anonymous):

I wrote the wrong slope and every other time I wrote the correct slope and the method was still correct. Normally you use what you wrote to solve for slope so I assumed you had misread the question. I agree truce sounds good.